A sample of a certain radioactive substance is prepared, and its initial activity is 4 µCi. After 2.4 hours, the activity is measured to be 1 µCi. What is the half-life of this material?
Answer #1

A=A0e-\lambdat

\lambda = 1/Ta =1/2.4=0.41

1=4*e-0.41t

log0.25=-0.41t

0.6/0.41=t

1.44=t